# Entanglement and apparent thermality in simulated black holes

### v1

@@ VI Discussion
We studied the mean-field theory limit of a chiral spin-chain simulator that, in its continuum limit, due to the emergence of an effective Lorentz invariance at low energies, describes a theory of Dirac fermions on a black hole space-time. For this model, we derived analytic expressions for the fermion zero-mode entropy of the black hole space-time’s interior and exterior regions. This entropy, which we derived under the assumption that the space-time is in thermal equilibrium with respect to any bipartition, is finite as a result of the spin-chain’s non-linear dispersion relation that naturally breaks the effective Lorentz invariance at the length scales of the lattice spacing.
By employing numerical simulations to investigate the spin-chain’s entanglement entropy and comparing it with our analytic black hole fermion zero-mode entropy, we identified regimes where the two entropies are equivalent and, by extension, where the assumption of thermality is valid and where it breaks down. Mainly, we demonstrated that, with respect to an arbitrary partition, the thermality condition is generally not valid. This is due to the absence of interactions, as the mean-field system cannot truly thermalise. Instead, somewhat surprisingly, the thermality condition only holds when the partition is located at the event horizon, due to the unique physics emerging in that region. We further verified this by studying the partitioned systems’ mode occupation, which revealed a thermal Fermi-Dirac distribution only for partitions near the event horizon.
@@ Acknowledgements
We are grateful to Cristian Voinea, Matthew Horner, Aiden Daniel, and Tanmay Bhore for their insightful discussions and guidance. This work was supported by EPSRC with Grant Nos. EP/W524372/1 and UKRI1337:Anyons24.
Appendix A. Chiral spin-chain continuum limit
In this appendix, we derive the low-energy and continuum limit of the chiral spin-chain Hamiltonian [24, 25, 26, 27], which describes a one-dimensional chain of $N$ interacting spin-1/2 spins and is defined as
$H=\sum^{N}_{n=1}\left[-\frac{u}{2}(\sigma^{x}_{n}\sigma^{x}_{n+1}+\sigma^{y}_{n}\sigma^{y}_{n+1})+\frac{v}{4}\chi_{n}\right],$ (A.1)
where $u,v\in\mathbb{R}$, $\{\sigma^{x}_{n},\sigma^{y}_{n},\sigma^{z}_{n}\}$ are the Pauli matrices acting on the $n$th spin, $\chi_{n}$ is a spin chirality operator [37, 38] given by
$\chi_{n}=\boldsymbol{\sigma}_{n}\cdot(\boldsymbol{\sigma}_{n+1}\crossproduct\boldsymbol{\sigma}_{n+2}),$ (A.2)
and $\boldsymbol{\sigma}_{n}=(\sigma^{x}_{n},\sigma^{y}_{n},\sigma^{z}_{n})$. We take periodic boundary conditions, such that $\boldsymbol{\sigma}_{N}=\boldsymbol{\sigma}_{0}$.
As demonstrated in Refs. [24, 25, 26, 27], due to the emergence of Lorentz invariance in its low-energy limit, the continuum limit of this Hamiltonian is described by the Hamiltonian of Dirac fermions in a curved space-time geometry. Demonstrating this begins with the application of a Jordan-Wigner transformation, defined as
$\begin{split}\sigma_{n}^{+}&=\left(\prod_{m<n}\left(1-2c^{\dagger}_{m}c_{m}\right)\right)c^{\dagger}_{n},\\ \sigma_{n}^{-}&=\left(\prod_{m<n}\left(1-2c^{\dagger}_{m}c_{m}\right)\right)c_{n},\end{split}$ (A.3)
where $c^{\dagger}_{n}$ and $c_{n}$ are fermionic creation and annihilation operators satisfying the anti-commutation relations $\{c_{n},c^{\dagger}_{m}\}=\delta_{nm}$ and $\{c_{n},c_{m}\}=\{c^{\dagger}_{n},c^{\dagger}_{m}\}=0$, and $\sigma^{z}_{n}=1-2c^{\dagger}_{n}c_{n}$. This Jordan-Wigner transformation maps the spins to fermions and the chiral spin-chain Hamiltonian of Eq. (A.1) to the following Hamiltonian of a system of interacting fermions: [24, 25, 26, 27]
$\begin{split}H=\sum^{N}_{n=1}\Biggl(&-uc^{\dagger}_{n}c_{n+1}-\frac{iv}{2}\Bigl(c^{\dagger}_{n}c_{n+2}-c^{\dagger}_{n}c_{n+1}\sigma^{z}_{n+2}\\ &\ \ \ \ \ \ -c^{\dagger}_{n+1}c_{n+2}\sigma^{z}_{n}\Bigr)\Biggr)+\text{H.c.}\end{split}$ (A.4)
The application of a mean-field theory approximation, valid in the regime of weak fluctuations, then allows the interacting fermion Hamiltonian of Eq. (A.4) to be described by the Hamiltonian [24, 25, 26, 27]
$H_{\text{MF}}=\sum^{N}_{n=1}\left(-uc^{\dagger}_{n}c_{n+1}-\frac{iv}{2}c^{\dagger}_{n}c_{n+2}\right)+\text{H.c.}$ (A.5)
Figure 6: The lattice corresponding to the chiral spin-chain model Hamiltonian of Eq. (1), where the lattice sites alternate between two sub-lattices, $A$ and $B$, and a two-site unit cell has been introduced.
By re-labelling the lattice sites such that they alternate between two sub-lattices, $A$ and $B$, and introducing a two-site unit cell, as illustrated in Fig. 6, the mean-field Hamiltonian of Eq. (A.5) can be rewritten as [24, 25, 26, 27]
$\begin{split}H_{MF}=\sum_{n}^{N}\biggl[&-ua^{\dagger}_{n}(b_{n}+b_{n-1})\\ &-\frac{iv}{2}(a^{\dagger}_{n}a_{n+1}+b^{\dagger}_{n}b_{n+1})\biggr]+\text{H.c.},\end{split}$ (A.6)
where $a^{\dagger}_{n},a_{n}$ and $b^{\dagger}_{n},b_{n}$ are fermionic creation and annihilation operators acting on the $n$th lattice sites of $A$ and $B$, respectively, which satisfy the fermionic anti-commutation relations $\{a_{n},a^{\dagger}_{m}\}=\{b_{n},b^{\dagger}_{m}\}=\delta_{nm}$ with all other anti-commutation relations being zero. Here, the index $n$ labels the two-site unit cell.
By performing the following Fourier transforms:
$a_{n}=\frac{1}{\sqrt{N_{c}}}\sum_{p}e^{ipa_{c}n}a_{p},\ \ \ b_{n}=\frac{1}{\sqrt{N_{c}}}\sum_{p}e^{ipa_{c}n}b_{p},$ (A.7)
where $p\in\{\pi,-\pi)$, $N_{c}=N/2$ is the number of two-sited unit cells, and $a_{c}$ is the lattice spacing of the two-site unit cell, the Hamiltonian of Eq. (A.6) can be expressed as [24, 25, 26, 27]
$H_{MF}=\sum_{p}\chi^{\dagger}_{p}h(p)\chi_{p},$ (A.8)
where $\chi_{p}=(a_{p},b_{p})^{T}$ is a two-component spinor and
$h(p)=\begin{pmatrix}g(p)&f(p)\\ f^{*}(p)&g(p)\end{pmatrix}$, (A.9)
$E(p)$ $=g(p)\pm|f(p)|,$ (A.10)
$=v\sin(a_{c}p)\pm u\sqrt{2+2\cos(a_{c}p)}.$ (A.11)
The dispersion relation of Eq. (A.11) has Fermi points located at momenta $p_{0}=\frac{\pi}{a_{c}}$ and $p_{\pm}=\pm\frac{1}{a_{c}}\arccos(1-\frac{2u^{2}}{v^{2}})$, of which only the former exists in the limit $v^{2}<u^{2}$ [24, 25, 26, 27].
The low-energy limit of the mean-field theory can be obtained by Taylor expanding the single-particle Hamiltonian of Eq. (A.9) and the dispersion relation of Eq. (A.11) around the Fermi point $p_{0}$ to $O(p^{2})$. Doing so gives
$h(p+p_{0})=-a_{c}v\mathbb{I}p+a_{c}u\sigma^{y}p\equiv e^{x}_{a}\alpha^{a}p,$ (A.12)
where the coefficients $e^{x}_{0}=-a_{c}v$ and $e^{x}_{1}=a_{c}u$ along with the matrices $\alpha^{0}=\mathbb{I}$ and $\alpha^{1}=\sigma^{y}$ have been defined, and $a\in\{0,1\}$. Following this, the continuum limit $a_{c}\rightarrow 0$ and the thermodynamics limit $N_{c}\rightarrow\infty$ can be taken, renormalising the couplings $a_{c}u\rightarrow u$ and $a_{c}v\rightarrow v$ and defining the coordinate $x=na_{c}$ in the process. Thus, in this continuum and low-energy limit, the mean-field Hamiltonian of Eq. (A.8) is given by [24, 25, 26, 27]
$H_{MF}=\int dp\chi^{\dagger}(p)e^{x}_{a}\alpha^{a}p\chi(p).$ (A.13)
Transforming the Hamiltonian of Eq. (A.13) into real space via the Fourier transformation
$\chi(p)=\frac{1}{\sqrt{L}}\int dxe^{-ipx}\chi(x)$ (A.14)
yields the following Hamiltonian: [24, 25, 26, 27]
$H_{MF}=\int dx\chi^{\dagger}(x)\left(-ie^{x}_{a}\alpha^{a}\overset{\leftrightarrow}{\partial}_{x}\right)\chi(x),$ (A.15)
where $\overset{\leftrightarrow}{\partial}_{\mu}=\frac{1}{2}(A\partial_{\mu}B-(\partial_{\mu}A)B)$. Lastly, following a Legendre transformation of the real space Hamiltonian of Eq. (A.15), the following action $S$ can be obtained:
$S$ $=i\int d^{1+1}x\chi^{\dagger}(x)\left(\overset{\leftrightarrow}{\partial}_{t}+e^{x}_{a}\gamma^{a}\overset{\leftrightarrow}{\partial}_{x}\right)\chi(x),$ (A.16)
$\equiv i\int d^{1+1}x\bar{\chi}(x)e^{\mu}_{a}\gamma^{a}\overset{\leftrightarrow}{\partial}_{\mu}\chi(x)$ (A.17)
where the coefficients $e^{t}_{0}=1$ and $e^{t}_{1}=0$, along with the flat space-time gamma matrices $\gamma^{a}=\sigma^{z}\alpha^{a}$ and the conjugate spinor $\bar{\chi}(x)=\chi^{\dagger}(x)\gamma^{0}$ have been defined, and $\mu\in\{t,x\}$ [24, 25, 26, 27]. Here, $|e|=\det(e^{a}_{\mu})=\sqrt{|g|}$, where $g$ is the determinant of the covariant metric tensor $g_{\mu\nu}$.
The action of Eq. (A.17) is analogous to that of Eq. (B.12) for a Dirac spinor $\psi$ on a (1+1)-dimensional Riemann-Cartan space-time, with the curved space-time spinor $\psi$ satisfying the anti-commutation relations of Eq. (B.20) and being related to the lattice spinor $\chi$ via the renormalisation $\chi=\sqrt{|e|}\psi$. The zweibein of the corresponding Riemann-Cartan space-time are given by [24, 25, 26, 27]
$e^{a}_{\mu}=\begin{pmatrix}1&v/u\\ 0&1/u\end{pmatrix},\ \ e^{\mu}_{a}=\begin{pmatrix}1&-v\\ 0&u\end{pmatrix}$, (A.18)
which correspond to the following covariant and contravariant metric tensors:
$g_{\mu\nu}=\begin{pmatrix}1-v^{2}/u^{2}&-v/u\\ -v/u&-1/u^{2}\end{pmatrix},\ \ g^{\mu\nu}=\begin{pmatrix}1&-v\\ -v&v^{2}-u^{2}\end{pmatrix}$, (A.19)
respectively, and the following metric line element:
$ds^{2}=\left(1-\frac{v^{2}}{u^{2}}\right)dt^{2}-\frac{2v}{u^{2}}dtdx-\frac{1}{u^{2}}dx^{2}.$ (A.20)
If the couplings $v$ and $u$ are slowly varying position-dependent functions $v(x)$ and $u(x)$, then the line element of Eq. (A.20) is that of the (1+1)-dimensional Gullstrand-Painlevé metric [39, 40] with the (+)(+\ -) metric signature [21, 24, 25, 26, 27], which describes the space-time geometry of a (1+1)-dimensional black hole with an event horizon of radius $x_{h}$ located at $v(x_{h})^{2}=u(x_{h})^{2}$ [21, 22], with the region bounded by $v(x)^{2}\geq u(x)^{2}$ corresponding the black hole’s interior [25, 24].
As the mean-field Hamiltonian of Eq. (A.5) is gapless for all $v(x)$, it can be described by a conformal field theory [25, 26]. It has been demonstrated [25, 26] that, for homogenous couplings, the regimes corresponding to the black hole space-time’s interior and exterior regions correspond to conformal field theories of central charges $c=2$ and $c=1$, respectively.
Appendix B. Spinor fields on a (1+1)D Riemann-Cartan space-time
@@ 1 Vielbein Formalism
Substitution of Eq. (B.3) for the relation between the vielbein and coordinate bases into Eq. (B.2) for the line element as expressed in the vielbein basis, then comparing to Eq. (B.1) for the line element as expressed in the coordinate basis, yields [57]
$g_{\mu\nu}(x)=\eta_{ab}e^{a}_{\mu}(x)e^{b}_{\nu}(x),$ (B.4)
@@
$e^{\mu}_{a}(x)e^{a}_{\nu}(x)=\delta^{\mu}_{\nu},\ \ \ \ e^{a}_{\mu}(x)e^{\mu}_{b}(x)=\delta^{a}_{b},$ (B.5)
which, along with Eqs. (B.3) and (B.4) for the definition of the vielbein and the general covariant metric tensor in terms of the vielbein, respectively, can be used to obtain
$g^{\mu\nu}(x)=\eta^{ab}e^{\mu}_{a}(x)e^{\nu}_{b}(x).$ (B.6)
@@ 2 Torsion, spin connection and Riemann-Cartan space-time
$\nabla_{\alpha}A^{\mu}_{\nu}=\partial_{\alpha}A^{\mu}_{\nu}+\Gamma^{\mu}_{\alpha\beta}A^{\beta}_{\nu}-\Gamma^{\beta}_{\alpha\nu}A^{\mu}_{\beta},$ (B.7)
@@
$\tilde{\Gamma}^{\alpha}_{\beta\gamma}=\frac{1}{2}g^{\alpha\mu}\left(\partial_{\gamma}g_{\beta\mu}+\partial_{\beta}g_{\gamma\mu}-\partial_{\mu}g_{\beta\gamma}\right).$ (B.8)
@@
$\omega^{a}_{\mu b}=e^{a}_{\alpha}\nabla_{\mu}e^{\alpha}_{b}=e^{a}_{\alpha}\left(\partial_{\mu}e^{\alpha}_{b}+\Gamma^{\alpha}_{\mu\beta}e^{\beta}_{b}\right),$ (B.9)
@@
$\nabla_{\mu}A^{a}_{b}=\partial_{\mu}A^{a}_{b}+\omega^{a}_{\mu c}A^{c}_{b}-\omega^{c}_{\mu b}A^{a}_{c}.$ (B.10)
@@ 3 Spinor field on a Riemann-Cartan space-time
$S_{RC}$ $=\int_{M}d^{1+1}x\mathcal{L}_{RC},$ (B.11)
$=\frac{i}{2}\int_{M}d^{1+1}x|e|\left(\bar{\psi}\gamma^{\mu}D_{\mu}\psi-\overline{D_{\mu}\psi}\gamma^{\mu}\psi+2im\bar{\psi}\psi\right),$ (B.12)
@@
$\omega_{\mu}=\frac{i}{2}\omega_{\mu ab}\Sigma^{ab},\ \ \ \Sigma^{ab}=\frac{i}{4}[\gamma^{a},\gamma^{b}],$ (B.15)
and $\omega_{\mu ab}=\eta_{ac}\omega^{c}_{\mu b}$. The notations $D_{\mu}$ and $\nabla_{\mu}$ are used to distinguish between the covariant derivative acting on spinors and tensors, respectively. From Eqs. (B.13) and (B.14) for the spinor covariant derivative and its adjoint, and the fact that $\{\omega_{\mu},\gamma^{\mu}\}\propto\{\gamma^{a},[\gamma^{b},\gamma^{c}]\}=0$ in (1+1)-dimensions, Eq. (B.12) for the action of a massive Dirac spinor on a (1+1)-dimensional Riemann-Cartan space-time can be expressed as
$S_{RC}=i\int_{M}d^{1+1}x|e|\left(\bar{\psi}\gamma^{\mu}\overset{\leftrightarrow}{\partial}_{\mu}\psi+2im\bar{\psi}\psi\right).$ (B.16)
@@
$\pi(t,x)=\frac{\partial\mathcal{L}}{\partial(\partial_{t}\psi)},$ (B.17)
@@
$\pi(t,x)=\frac{i}{2}|e|\psi^{\dagger}(t,x)\gamma^{0}\gamma^{t}.$ (B.18)
To quantise the action for the Dirac spinor on a general Riemann-Cartan space-time given in Eq. (B.12), $\psi(t,x)$ and $\pi(t,x)$ are promoted to Hermitian operators $\hat{\psi}(t,x)$ and $\hat{\pi}(t,x)$ and are required to satisfy the following equal-time canonical anti-commutation relations [57]:
$\begin{split}\{\hat{\psi}_{\alpha}(t,x),\hat{\psi}_{\beta}(t,x^{\prime})\}=\{\hat{\pi}&{}_{\alpha}(t,x),\hat{\pi}_{\beta}(t,x^{\prime})\}=0,\\ \{\hat{\psi}_{\alpha}(t,x),\hat{\pi}_{\beta}(t,x^{\prime})\}&=i\delta_{\alpha\beta}\delta(x-x^{\prime}),\end{split}$ (B.19)
where $\alpha,\beta$ denote spinor components. Using these canonical anti-commutation relations and Eq. (B.18) for the canonical momentum of a Dirac spinor, one obtains the following additional anti-commutation relation between the field operator $\hat{\psi}$ and its Hermitian conjugate $\hat{\psi}^{\dagger}$:
$\{\hat{\psi}_{\alpha}(t,x),\hat{\psi}_{\beta}^{\dagger}(t,x)\}=\frac{1}{|e|}(\gamma^{0}\gamma^{t})^{-1}\delta_{\alpha\beta}\delta(x-x^{\prime}).$ (B.20)
If the space-time $M$ is described by a stationary metric (that is, a metric that is invariant under the time-translation transformation $t\mapsto t+t_{0}$, where $t_{0}\in\mathbb{R}$), then it will admit a time-like Killing vector $\xi^{\mu}=\delta^{\mu}_{t}$ (often equivalently denoted $\partial_{t}$) satisfying the Killing equation of Eq (C.1). Consequently, the energy of the spinor will be a well-defined quantity that is conserved along the metric’s geodesics [57, 23]. This allows a Hamiltonian $H$ to be defined via the following Legendre transformation of the Lagrangian density: [59]
$H=\int_{S}dx\left(\frac{\partial\mathcal{L}}{\partial(\partial_{t}\psi)}\partial_{t}\psi-\mathcal{L}\right),$ (B.21)
@@
Appendix C. Hawking effect in chiral black hole metric

### v2

@@ VI Discussion
We studied the mean-field theory limit of a chiral spin-chain simulator that, in its continuum limit, due to the emergence of an effective Lorentz invariance at low energies, describes a theory of Dirac fermions on a black hole space-time. For this model, we obtained analytic expressions for the fermion zero-mode entropy of the black hole space-time’s interior and exterior regions. This entropy, derived under the assumption that the fermions are suitably described by a thermal Gibbs state with respect to any space-time bipartition, is finite as a result of the spin-chain’s non-linear dispersion relation, which naturally breaks the effective Lorentz invariance at the length scales of the lattice spacing.
By employing numerical simulations to investigate the spin-chain’s entanglement entropy and comparing it with our analytic black hole fermion zero-mode entropy, we identified regimes where the two entropies are equivalent and, by extension, where the assumption of thermality is valid and where it breaks down. Mainly, we demonstrated that, with respect to an arbitrary partition, the thermality condition is generally not valid. This is due to the absence of interactions, as the mean-field system cannot truly thermalise. Instead, somewhat surprisingly, the thermality condition holds only when the partition is located at the event horizon, due to the unique physics emerging in that region. We further verified this by studying the mode occupation of the partitioned system, which revealed a thermal Fermi-Dirac distribution only for partitions near the event horizon.
@@ Acknowledgements
We are grateful to Cristian Voinea, Matthew Horner, Aiden Daniel, and Tanmay Bhore for their insightful discussions and guidance. This work was supported by EPSRC with Grant Nos. EP/W524372/1 and UKRI1337:Anyons24. Statement of compliance with EPSRC policy framework on research data: This publication is theoretical work that does not require supporting research data. A.H. acknowledges support from Leverhulme Trust Research Leadership Award RL-2019-015.
Appendix A. Spinor fields on a (1+1)D Riemann-Cartan space-time
@@ 1 Vielbein Formalism
Substitution of Eq. (A.3) for the relation between the vielbein and coordinate bases into Eq. (A.2) for the line element as expressed in the vielbein basis, then comparing to Eq. (A.1) for the line element as expressed in the coordinate basis, yields [48]
$g_{\mu\nu}(x)=\eta_{ab}e^{a}_{\mu}(x)e^{b}_{\nu}(x),$ (A.4)
@@
$e^{\mu}_{a}(x)e^{a}_{\nu}(x)=\delta^{\mu}_{\nu},\ \ \ \ e^{a}_{\mu}(x)e^{\mu}_{b}(x)=\delta^{a}_{b},$ (A.5)
which, along with Eqs. (A.3) and (A.4) for the definition of the vielbein and the general covariant metric tensor in terms of the vielbein, respectively, can be used to obtain
$g^{\mu\nu}(x)=\eta^{ab}e^{\mu}_{a}(x)e^{\nu}_{b}(x).$ (A.6)
@@ 2 Torsion, spin connection and Riemann-Cartan space-time
$\nabla_{\alpha}A^{\mu}_{\nu}=\partial_{\alpha}A^{\mu}_{\nu}+\Gamma^{\mu}_{\alpha\beta}A^{\beta}_{\nu}-\Gamma^{\beta}_{\alpha\nu}A^{\mu}_{\beta},$ (A.7)
@@
$\tilde{\Gamma}^{\alpha}_{\beta\gamma}=\frac{1}{2}g^{\alpha\mu}\left(\partial_{\gamma}g_{\beta\mu}+\partial_{\beta}g_{\gamma\mu}-\partial_{\mu}g_{\beta\gamma}\right).$ (A.8)
@@
$\omega^{a}_{\mu b}=e^{a}_{\alpha}\nabla_{\mu}e^{\alpha}_{b}=e^{a}_{\alpha}\left(\partial_{\mu}e^{\alpha}_{b}+\Gamma^{\alpha}_{\mu\beta}e^{\beta}_{b}\right),$ (A.9)
@@
$\nabla_{\mu}A^{a}_{b}=\partial_{\mu}A^{a}_{b}+\omega^{a}_{\mu c}A^{c}_{b}-\omega^{c}_{\mu b}A^{a}_{c}.$ (A.10)
@@ 3 Spinor field on a Riemann-Cartan space-time
$S_{RC}$ $=\int_{M}d^{1+1}x\mathcal{L}_{RC},$ (A.11)
$=\frac{i}{2}\int_{M}d^{1+1}x|e|\left(\bar{\psi}\gamma^{\mu}D_{\mu}\psi-\overline{D_{\mu}\psi}\gamma^{\mu}\psi+2im\bar{\psi}\psi\right),$ (A.12)
@@
$\omega_{\mu}=\frac{i}{2}\omega_{\mu ab}\Sigma^{ab},\ \ \ \Sigma^{ab}=\frac{i}{4}[\gamma^{a},\gamma^{b}],$ (A.15)
and $\omega_{\mu ab}=\eta_{ac}\omega^{c}_{\mu b}$. The notations $D_{\mu}$ and $\nabla_{\mu}$ are used to distinguish between the covariant derivative acting on spinors and tensors, respectively.
From Eqs. (A.13) and (A.14) for the spinor covariant derivative and its adjoint, and the fact that $\{\omega_{\mu},\gamma^{\mu}\}\propto\{\gamma^{a},[\gamma^{b},\gamma^{c}]\}=0$ in (1+1)-dimensions, Eq. (A.12) for the action of a massive Dirac spinor on a (1+1)-dimensional Riemann-Cartan space-time can be expressed as
$S_{RC}=i\int_{M}d^{1+1}x|e|\left(\bar{\psi}\gamma^{\mu}\overset{\leftrightarrow}{\partial}_{\mu}\psi+im\bar{\psi}\psi\right).$ (A.16)
@@
$\pi(t,x)=\frac{\partial\mathcal{L}}{\partial(\partial_{t}\psi)},$ (A.17)
@@
$\pi(t,x)=\frac{i}{2}|e|\psi^{\dagger}(t,x)\gamma^{0}\gamma^{t}.$ (A.18)
To quantise the action for the Dirac spinor on a general Riemann-Cartan space-time given in Eq. (A.12), $\psi(t,x)$ and $\pi(t,x)$ are promoted to Hermitian operators $\hat{\psi}(t,x)$ and $\hat{\pi}(t,x)$ and are required to satisfy the following equal-time canonical anti-commutation relations [48]:
$\begin{split}\{\hat{\psi}_{\alpha}(t,x),\hat{\psi}_{\beta}(t,x^{\prime})\}=\{\hat{\pi}&{}_{\alpha}(t,x),\hat{\pi}_{\beta}(t,x^{\prime})\}=0,\\ \{\hat{\psi}_{\alpha}(t,x),\hat{\pi}_{\beta}(t,x^{\prime})\}&=i\delta_{\alpha\beta}\delta(x-x^{\prime}),\end{split}$ (A.19)
where $\alpha,\beta$ denote spinor components. Using these canonical anti-commutation relations and Eq. (A.18) for the canonical momentum of a Dirac spinor, one obtains the following additional anti-commutation relation between the field operator $\hat{\psi}$ and its Hermitian conjugate $\hat{\psi}^{\dagger}$:
$\{\hat{\psi}_{\alpha}(t,x),\hat{\psi}_{\beta}^{\dagger}(t,x)\}=\frac{1}{|e|}(\gamma^{0}\gamma^{t})^{-1}\delta_{\alpha\beta}\delta(x-x^{\prime}).$ (A.20)
If the space-time $M$ is described by a stationary metric (that is, a metric that is invariant under the time-translation transformation $t\mapsto t+t_{0}$, where $t_{0}\in\mathbb{R}$), then it will admit a time-like Killing vector $\xi^{\mu}=\delta^{\mu}_{t}$ (often equivalently denoted $\partial_{t}$) satisfying the Killing equation of Eq (B.1). Consequently, the energy of the spinor will be a well-defined quantity that is conserved along the metric’s geodesics [48, 35]. This allows a Hamiltonian $H$ to be defined via the following Legendre transformation of the Lagrangian density: [21]
$H=\int_{S}dx\left(\frac{\partial\mathcal{L}}{\partial(\partial_{t}\psi)}\partial_{t}\psi-\mathcal{L}\right),$ (A.21)
@@
Appendix B. Hawking effect in chiral black hole metric